# MonSter: Marry Monodepth to Stereo Unleashes Power

### v1

@@ 3 Method
Figure 4: Overview of MonSter. MonSter consists of a monocular depth estimation branch, a stereo matching branch, and a mutual refinement module. It iteratively improves one branch with priors from the other, effectively resolving the ill-posedness in stereo matching.
@@
As shown in Fig. 4, MonSter consists of 1) a monocular depth branch, 2) a stereo matching branch and 3) a mutual refinement module. The two branches estimate initial monocular depth and the stereo disparity, which are fed into mutual refinement to iteratively improve each other.
3.1 Monocular and Stereo Branches
@@
The monocular depth branch can leverage most monocular depth models to achieve non-trivial performance improvements (see Sec.4.5 for analysis). Our best empirical configuration uses pretrained DepthAnythingV2 [46] as the monocular depth branch, which uses DINOv2 [24] as the encoder and DPT [26] as the decoder. The stereo matching branch follows IGEV [38] to obtain the initial stereo disp, with modifications only to the feature extraction component, as shown in Fig. 4. To efficiently and fully leverage the pretrained monocular model, the stereo branch shares the ViT encoder in DINOv2 with the monocular branch, with parameters frozen to prevent the stereo matching training from affecting its generalization ability. Moreover, the ViT architecture extracts feature at a single resolution, while recent stereo matching methods commonly utilize multi-scale features at four scales (1/32, 1/16, 1/8, and 1/4 of the original image resolution ). To fully align with IGEV, we employ a stack of 2D convolutional layers, denoted as the feature transfer network, to downsample and transform the ViT feature into a collection of pyramid features =$\mathcal{F}=\{\bm{F}_{0},\bm{F}_{1},\bm{F}_{2},\bm{F}_{3}\}$, where $\bm{F}_{k}\in\mathbb{R}^{\frac{H}{2^{5-k}}\times\frac{W}{2^{5-k}}\times{c}_{k}}$. We follow IGEV to construct Geometry Encoding Volume and use the same ConvGRUs for iterative optimization. To balance accuracy and efficiency, we perform only $N_{1}$ iterations to obtain a initial stereo disparity with reasonable quality.
3.2 Mutual Refinement
@@
where $\bm{F}_{3}^{L}$,$\bm{F}_{3}^{R}$ represent the quarter-resolution features of left and right images in $\mathcal{F}$ respectively. For each iteration, we also use the current stereo disparity $\bm{D}_{S}^{j}$ to index from the Geometry Encoding Volume to obtain geometry features of stereo branch $\bm{G}_{S}^{j}$ follow IGEV. Concatenated with $\bm{F}_{S}^{j}$ and $\bm{D}_{S}^{j}$, we obtain the stereo condition as:
@@
Figure 5: Qualitative results on ETH3D. MonSter outperforms IGEV in challenging areas with strong reflectance, fine structures etc.
@@
where $\bm{G}_{M}^{j}$ is the geometry features of the monocular branch obtained by indexing from the Geometry Encoding Volume using the monocular disparity $\bm{D}_{M}^{j}$. $x_{M}^{j}$ represents the monocular condition for ConvGRUs, and we use the Eq. (4) to update the hidden state $h_{S}^{i}$ similarly, with only the condition input changed from $x_{S}^{i}$ to $x_{M}^{i}$. We use the same two convolutional layers to decode the residual disparity $\triangle\bm{d}$ and update the current stereo disparity following Eq. (5). After $N_{2}$ rounds of dual-branched refinement, the disparity of the stereo branch is the final output of MonSter.
3.3 Loss Function
@@
$\mathcal{L}_{Stereo}=$ $\;\sum_{i=0}^{{N}_{1}-1}\gamma^{{N}_{1}+{N}_{2}-i}||\bm{d}_{i}-\bm{d}_{gt}||_{1}+$ (7)
$\;\sum_{i={N}_{1}}^{{N}_{1}+{N}_{2}-1}\gamma^{{N}_{1}+{N}_{2}-i}||\bm{D}_{S}^{i-{N}_{1}}-\bm{d}_{gt}||_{1},$
$\mathcal{L}_{Mono}=$ $\;\sum_{i={N}_{1}}^{{N}_{1}+{N}_{2}-1}\gamma^{{N}_{1}+{N}_{2}-i}||\bm{D}_{m}^{i-{N}_{1}}-\bm{d}_{gt}||_{1}$
@@
4 Experiment
Method GwcNet[14] LEAStereo[8] ACVNet[37] IGEV[38] Selective-IGEV[35] MonSter (Ours)
EPE (px)$\downarrow$ 0.76 0.78 0.48 0.47 0.44 0.37 (-15.91%)
Table 1: Quantitative evaluation on Scene Flow test set. The best result is bolded, and the second-best result is underscored.
@@ 4 Experiment
LEAStereo[8] - - - 7.15 12.10 8.11 1.29 1.51 1.40 1.65 1.90 2.39 1.13 1.45
ACVNet[37] 2.58 2.86 0.45 13.70 19.50 32.2 1.26 1.52 1.37 1.65 1.83 2.35 1.13 1.47
RAFT-Stereo[20] 2.44 2.60 0.36 4.74 9.37 8.41 1.44 1.69 1.58 1.82 1.92 2.42 1.30 1.66
CREStereo[16] 0.98 1.09 0.28 3.71 8.13 7.70 1.33 1.54 1.45 1.69 1.72 2.18 1.14 1.46
IGEV[38] 1.12 1.51 0.34 4.83 8.16 12.80 1.27 1.49 1.38 1.59 1.71 2.17 1.12 1.44
CroCo-Stereo[36] 0.99 1.14 0.30 7.29 11.11 8.91 1.30 1.51 1.38 1.59 - - - -
DLNR[49] - - - 3.20 6.98 7.78 1.42 1.61 1.60 1.76 - - - -
Selective-IGEV[35] 1.23 1.56 0.29 2.51 6.04 7.26 1.22 1.44 1.33 1.55 1.59 2.05 1.07 1.38
LoS[17] 0.91 1.03 0.31 4.20 8.03 6.99 1.29 1.52 1.42 1.65 1.69 2.12 1.10 1.38
NMRF-Stereo[13] - - - - - - 1.18 1.46 1.28 1.57 1.59 2.07 1.01 1.35
MonSter(Ours) 0.46 0.72 0.20 2.64 6.14 6.71 1.05 1.33 1.13 1.41 1.36 1.75 0.84 1.09
Table 2: Results on four popular benchmarks. All results are derived from official leaderboard publications or corresponding papers. All metrics are presented in percentages, except for RMSE, which is reported in pixels. For testing masks, “All” denotes being tested with all pixels while “Noc” denotes being tested with a non-occlusion mask. The best and second best are marked with colors.
4.1 Implementation Details
@@
We implement MonSter with PyTorch and perform experiments using NVIDIA RTX 3090 GPUs. We use the AdamW [21] optimizer and clip gradients to the range [-1, 1] following baseline [38]. We use the one-cycle learning rate schedule with a learning rate of 2e-4 and train MonSter with a batch size of 8 for 200k steps as the pretrained model. For the monocular branch, we use the ViT-large version of DepthAnythingV2 [46] and freeze its parameters to prevent training of stereo-matching tasks from affecting its generalization and accuracy.
Following the standard [16, 17, 35], we pretrain MonSter on Scene Flow [22] for most experiments. For finetuning on ETH3D and Middlebury, we follow SOTA methods [16, 17, 35] to create the Basic Training Set (BTS) from various public datasets for pretraining, including Scene Flow [22], CREStereo [16], Tartan Air [34], Sintel Stereo [3], FallingThings [32] and InStereo2k [2].
4.2 Benchmark Performance
@@
To demonstrate the outstanding performance of our method, we evaluate MonSter on the five most commonly used benchmarks: KITTI 2012 [11], KITTI 2015 [23], ETH3D [28], Middlebury [27] and Scene Flow [22].
@@
Scene Flow [22]. As shown in Tab. 1, we achieve a new state-of-the-art performance with an EPE metric of 0.37 on Scene Flow, surpassing our baseline [38] by 21.28% and outperforming the SOTA method [35] by 15.91%.
@@
ETH3D[28]. Following the SOTA methods [16, 17, 40, 42, 35], the full training set is composed of the BTS and ETH3D training set. Our MonSter ranks 1st on the ETH3D leaderboard. As shown in Tab. 2, MonSter achieves the best performance among all published methods for all metrics. Compared with baseline IGEV, we achieve improvements of 58.93%, 52.32%, and 41.18% in the three reported metrics, respectively. Qualitative comparisons shown in Fig.5 exhibit a similar trend. Notably, even compared with the previous best method LoS, we improved the Bad 1.0 (Noc) metric from 0.91 to 0.46, achieving a 49.45% improvement.
@@
Middlebury. Also following  [16, 17, 42, 35], the training set of Middlebury is the combination of the BTS and Middlebury training set. As shown in Tab. 2, MonSter outperforms all existing methods in terms of RMSE metric. Compared with baseline [38], we achieved an improvement of 45.34% on the Bad 2.0 (Noc) metric.
@@
KITTI. Following the training of SOTA methods  [16, 13, 35], we finetuned Scene Flow pretrained model on the mixed dataset of KITTI 2012 and KITTI 2015 for 50k steps. At the time of writing, MonSter ranks $1^{st}$ simultaneously on both the KITTI 2012 and KITTI 2015 leaderboards. As shown in Tab. 2, we achieve the best performance for all metrics. On KITTI 2015, MonSter surpasses CREStereo [16] and Selective-IGEV [35] by 16.57% and 9.03% on D1-all metric of all regions, respectively. As for KITTI 2012, we significantly outperform the existing SOTA by a large margin. Compared to the current highest-precision method, NMRF-Stereo [13] and Selective-IGEV [35], we achieve a 19.26% improvement and a 21.01% improvement respectively in the Out-3 metric across all regions.
4.3 Performance in Ill-posed Regions

### v2

@@ II-D Multi-view depth estimation with Structural Priors
Figure 5: Overview of MonSter++. MonSter++ consists of a monocular depth estimation branch, a stereo matching branch, and a mutual refinement module. It iteratively improves one branch with priors from the other, effectively resolving the ill-posedness in stereo matching.
III Method
We propose MonSter++, a unified framework that can be applied to both stereo matching and multi-view stereo for accurate metric depth estimation. For both tasks, the network architecture remains identical, with the only difference in the cost volume construction—a parameter-free process that operates by directly measuring feature similarity. Therefore, we first detail the MonSter++ framework using stereo matching as an example (Sec. III-A), then explicitly contrast it with the multi-view stereo model (Sec. III-B). In addition, we describe details of our real-time version of MonSter++, named RT-MonSter++ (Sec. III-C).
III-A Stereo matching
@@
As shown in Fig. 5, MonSter++ consists of 1) a monocular depth branch, 2) a stereo matching branch, and 3) a mutual refinement module. The two branches estimate initial monocular depth and the stereo disparity, which are fed into mutual refinement to iteratively improve each other.
III-A1 Monocular and Stereo Branches
@@
The monocular depth branch can leverage most monocular depth models to achieve non-trivial performance improvements (see Sec.IV-G for analysis). Our best empirical configuration uses pretrained DepthAnythingV2 [38] as the monocular depth branch, which uses DINOv2 [69] as the encoder and DPT [88] as the decoder. The stereo matching branch follows IGEV [6] to obtain the initial stereo disparity, with modifications only to the feature extraction component, as shown in Fig. 5. To efficiently and fully leverage the pretrained monocular model, the stereo branch shares the ViT encoder in DINOv2 with the monocular branch, with parameters frozen to prevent the stereo matching training from affecting its generalization ability. Moreover, the ViT architecture extracts feature at a single resolution, while recent stereo matching methods commonly utilize multi-scale features at four scales (1/32, 1/16, 1/8, and 1/4 of the original image resolution). To fully align with IGEV, we employ a stack of 2D convolutional layers, denoted as the feature transfer network, to downsample and transform the ViT feature into a collection of pyramid features $\mathcal{F}=\{\bm{F}_{0},\bm{F}_{1},\bm{F}_{2},\bm{F}_{3}\}$, where $\bm{F}_{k}\in\mathbb{R}^{\frac{H}{2^{5-k}}\times\frac{W}{2^{5-k}}\times{c}_{k}}$. We follow IGEV to construct the Geometry Encoding Volume and use the same ConvGRUs for iterative optimization. To balance accuracy and efficiency, we perform only $N_{1}$ iterations to obtain an initial stereo disparity with reasonable quality.
III-A2 Mutual Refinement
@@
where $\bm{F}_{3}^{L}$$\bm{F}_{3}^{R}$ represent the quarter-resolution features of left and right images in $\mathcal{F}$ respectively. For each iteration, we also use the current stereo disparity $\bm{D}_{S}^{j}$ to index from the Geometry Encoding Volume to obtain geometry features of stereo branch $\bm{G}_{S}^{j}$ follow IGEV. Concatenated with $\bm{F}_{S}^{j}$ and $\bm{D}_{S}^{j}$, we obtain the stereo condition as:
@@
where $\bm{G}_{M}^{j}$ is the geometry features of the monocular branch obtained by indexing from the Geometry Encoding Volume using the monocular disparity $\bm{D}_{M}^{j}$. $x_{M}^{j}$ represents the monocular condition for ConvGRUs, and we use Eq. (4) to update the hidden state $h_{S}^{i}$ similarly, with only the condition input changed from $x_{S}^{i}$ to $x_{M}^{i}$. We use the same two convolutional layers to decode the residual disparity $\triangle\bm{d}$ and update the current stereo disparity following Eq. (5). After $N_{2}$ rounds of dual-branched refinement, the disparity of the stereo branch is the final output of MonSter++.
III-A3 Loss Function
@@
[…]
@@
Figure 6: Overview of RT-MonSter++, the real-time variant of MonSter++, capable of achieving over 20 FPS inference at 1K resolution.
III-B Multi-view stereo
The fundamental difference between multi-view stereo and stereo matching lies in their input: a set of unrectified images with arbitrary camera poses, where matching must be performed along more general epipolar geometries. This difference leads to a key distinction in the cost volume construction stage.
Stereo matching only requires cost computation along horizontal directions between two views, we adopt the Geometry Encoding Volume following IGEV [6] in our stereo model, as described in Sec. III-A1. In multi-view stereo setting, we follow MVSNet [8] to construct a variance-based cost metric that encodes feature similarity across multiple unrectified views to replace the Geometry Encoding Volume. For simplicity, in the following we denote $\mathbf{I}_{1}$ as the reference image, $\{\mathbf{I}_{i}\}_{i=2}^{N}$ the source images, and $\{\mathbf{K}_{i},\mathbf{R}_{i},\mathbf{t}_{i}\}_{i=1}^{N}$ the camera intrinsics, rotations and translations that correspond to the feature maps. Following MVSNet [8], we wrap features from source views to the reference view using differentiable homography:
$\mathbf{H}_{i}(d)=\mathbf{K}_{i}\cdot\mathbf{R}_{i}\cdot\Big(\mathbf{I}-\frac{(\mathbf{t}_{1}-\mathbf{t}_{i})\cdot\mathbf{n}_{1}^{T}}{d}\Big)\cdot\mathbf{R}_{1}^{T}\cdot\mathbf{K}_{1}^{T}.$ (8)
where $d\in\mathbb{B}$, $\mathbb{B}$ denotes the bins uniformly sampled in the log space from $d_{min}$ to $d_{max}$, which represents the depth search range. Then we use variance-based cost metric $\mathcal{M}$ for N-view similarity measurement. We aggregate multiple feature volumes $\{\mathbf{V}_{i}\}_{i=1}^{N}$ to one cost volume $\mathbf{C}$ as:
$\mathbf{C}=\mathcal{M}(\mathbf{V}_{1},\cdots,\mathbf{V}_{N})=\frac{\sum\limits_{i=1}^{N}{(\mathbf{V}_{i}-{\mathbf{V}_{i}})^{2}}}{N}$ (9)
The variance-based cost volume $\mathbf{C}$ shares the same dimensions as the Geometry Encoding Volume used in the stereo matching model, and is entirely parameter-free. The subsequent processing stages remain identical to those in the stereo matching pipeline.
III-C Real-time version of MonSter++
Building upon MonSter++, we develop a real-time variant capable of achieving over 20 FPS inference speed at 1K resolution with only 2 GB of GPU memory. While MonSter++ performs cost aggregation and iterative refinement solely at the $1/4$ resolution level, our real-time model adopts a fully coarse-to-fine framework, progressively operating at $1/16$, $1/8$, and $1/4$ resolutions as shown in Fig.6. This hierarchical design enables early-stage pruning of the disparity search space, significantly improving efficiency. To further reduce computational overhead at higher resolutions ($1/8$ and $1/4$), we construct local cost volumes instead of full ones. This cascaded search and multi-scale depth fusion strategy enables us to fully exploit monocular priors while achieving both high efficiency and strong robustness. Specifically, for each pixel, we center the disparity sampling around the estimate from the previous resolution stage. Let $d_{\text{prev}}(i)$ denote the predicted disparity in pixel $i$ from the preceding stage, and $W$ denote the feature width at the current stage. Given the number of disparity samples $D$ and the sampling interval $\delta$, the local disparity range is defined as follows:
$d_{\min}(i)$ $=\max\left(0,\,d_{\text{prev}}(i)-\tfrac{D-1}{2}\cdot\delta\right),$ (10)
$d_{\max}(i)$ $=\min\left(W,\,d_{\text{prev}}(i)+\tfrac{D-1}{2}\cdot\delta\right).$
To account for boundary clipping, we define the actual sampling interval $\Delta$ as:
$\Delta(i)=\frac{d_{\max}(i)-d_{\min}(i)}{D-1}.$ (11)
Finally, the set of sampled disparities at that pixel is given by:
$d_{\text{sample}}(i)=\left\{d_{\min}(i)+n\cdot\Delta(i)\mid n=0,1,\ldots,D-1\right\}.$ (12)
This local sampling strategy effectively narrows the disparity search range at each level, leading to a substantial speed-up without compromising accuracy.
To improve runtime efficiency while maintaining competitive accuracy, we make several adjustments to the model architecture. First, our real-time lightweight variant adopts a single-layer ConvGRU. Since our design follows a coarse-to-fine framework, only the hidden state corresponding to the current resolution stage needs to be updated, which further improves efficiency without sacrificing accuracy. Second, we reduce the number of GRU updates: one update at each of the two lower resolutions and two updates at 1/4 resolution. Finally, we redesign the cost aggregation module using a lightweight architecture to better balance speed and performance.
IV Experiment
TABLE I: Quantitative evaluation on Scene Flow test set. The best result is bolded, and the second-best result is underscored.
Method GwcNet[2] LEAStereo[89] ACVNet[26] IGEV[6] Selective-IGEV[5] MonSter++ (Ours)
[…]
TABLE II: Results on four popular stereo matching benchmarks. All results are derived from official leaderboard publications or corresponding papers. All metrics are presented in percentages, except for RMSE, which is reported in pixels. For testing masks, “All” denotes being tested with all pixels while “Noc” denotes being tested with a non-occlusion mask. The best and second best are marked with colors.
@@ IV Experiment
LEAStereo[89] - - - 2.75 8.11 13.7 1.29 1.51 1.40 1.65 1.90 2.39 1.13 1.45
ACVNet[26] 2.58 2.86 0.45 7.99 32.2 38.6 1.26 1.52 1.37 1.65 1.83 2.35 1.13 1.47
RAFT-Stereo[4] 2.44 2.60 0.36 2.75 8.41 12.6 1.44 1.69 1.58 1.82 1.92 2.42 1.30 1.66
CREStereo[30] 0.98 1.09 0.28 2.04 7.70 10.5 1.33 1.54 1.45 1.69 1.72 2.18 1.14 1.46
IGEV[6] 1.12 1.51 0.34 3.33 12.80 15.1 1.27 1.49 1.38 1.59 1.71 2.17 1.12 1.44
CroCo-Stereo[90] 0.99 1.14 0.30 4.18 8.91 10.6 1.30 1.51 1.38 1.59 - - - -
DLNR[31] - - - 1.89 7.78 10.2 1.42 1.61 1.60 1.76 - - - -
Selective-IGEV[5] 1.23 1.56 0.29 1.36 7.26 9.26 1.22 1.44 1.33 1.55 1.59 2.05 1.07 1.38
LoS[86] 0.91 1.03 0.31 2.30 6.99 8.78 1.29 1.52 1.42 1.65 1.69 2.12 1.10 1.38
[…]
IV-A Implementation
@@
We implement MonSter++ with PyTorch and perform experiments using NVIDIA RTX 4090 GPUs. We use the AdamW [92] optimizer and clip gradients to the range [-1, 1] following baseline [6]. We use the one-cycle learning rate schedule with a learning rate of 2e-4 and train MonSter++ with a batch size of 8 for 200k steps as the pretrained model. For the monocular branch, we use the ViT-large version of DepthAnythingV2 [38] and freeze its parameters to prevent training of stereo-matching tasks from affecting its generalization and accuracy  [^1] All research undertaken at Meta AI was limited to general guidance on model architectural design. Meta did not participate in any model training activities..
Following the standard [30, 86, 5, 87], we adopt a two-stage training strategy for the evaluation on the ETH3D and Middlebury benchmarks. In the first training stage, we follow SOTA methods [30, 86, 5, 87] to create the Basic Training Set (BTS) from various public datasets for pretraining, including Scene Flow [22], FSD [87], CREStereo [30], Tartan Air [93], Sintel Stereo [94], FallingThings [95], and InStereo2k [96]. To develop a highly generalizable stereo model for open-source community use, we additionally collected and aggregated a wide range of public datasets, resulting in over 2 million image pairs for training. We refer to this combined dataset as the Full Training Set (FTS). Compared to the Basic Training Set (BTS), the Full Training Set (FTS) additionally includes data from 3DKenBurns [97], DynamicStereo [98], IRS [99], VA [100], Booster [101], and Carla-Highres [102] datasets.
IV-B Stereo matching Benchmark Performance
@@
To demonstrate the outstanding performance of our method, we independently evaluate MonSter++ on both stereo matching and multi-view stereo tasks. We first evaluate our method on the stereo matching task using the five most commonly used benchmarks: KITTI 2012 [39], KITTI 2015 [23], ETH3D [20], Middlebury [21], and Scene Flow [22].
@@
Scene Flow [22]. As shown in Tab. I, we achieve a new state-of-the-art performance with an EPE metric of 0.37 on Scene Flow, surpassing our baseline [6] by 21.28% and outperforming the SOTA method [5] by 15.91%.
ETH3D[20]. Following the SOTA methods [30, 86, 33, 29, 5], the full training set is composed of the BTS and ETH3D training set.
@@
Our MonSter++ ranks 1st on the ETH3D leaderboard. As shown in Tab. II, MonSter++ achieves the best performance among all published methods. Compared with baseline IGEV, we achieve improvements of 77.68%, 70.20%, and 47.06% in the three reported metrics, respectively. Qualitative comparisons shown in Fig.7 exhibit a similar trend. Notably, even compared with the previous best method FoundationStereo, we improved the RMSE (Noc) metric from 0.20 to 0.18, achieving a 10% improvement.
TABLE III: Results on two popular multi-view stereo depth estimation benchmarks. Note that the * marks the result reproduced by us using their open-source code, other reported numbers are from  [13] and the corresponding original papers. The best and second best are marked with colors.
Method DDAD [24] KITTI Eigen Split [25]
[…]
Deepv2d [103] - - - 0.091 0.582 3.644 0.154 0.923
CasMVS [9] 0.129 2.01 9.87 0.066 0.228 2.567 0.112 0.945
MVSNet [8] 0.109 1.62 8.21 - - - - -
IterMVS [10] 0.104 1.59 7.95 0.057 0.178 2.234 0.087 0.968
MVS2D [11] 0.132 2.05 9.82 0.058 0.176 2.277 0.090 0.966
SC-GAN [104] - - - 0.063 0.178 2.129 0.097 0.961
MaGNet [85] 0.112 1.74 9.23 0.054 0.162 2.158 0.083 0.971
[…]
Figure 7: Qualitative results on ETH3D. MonSter++ outperforms IGEV in challenging areas with strong reflectance, fine structures etc.
@@
Middlebury. Also following  [30, 86, 29, 5], the training set of Middlebury is the combination of the BTS and Middlebury training set. As shown in Tab. II, MonSter++ outperforms all existing methods in terms of the RMSE metric. Compared with baseline [6], we achieve an improvement of 64.56% on the Bad 4.0 (Noc) metric. Compared with the SOTA method FoundationStereo [87], we achieve a 6.94% improvement in the RMSE (Noc) metric.
@@
KITTI. Following the training of SOTA methods  [30, 91, 5], we finetuned the Scene Flow pretrained model on the mixed dataset of KITTI 2012 and KITTI 2015 for 50k steps. At the time of writing, MonSter++ ranks $1^{st}$ simultaneously on both the KITTI 2012 and KITTI 2015 leaderboards. As shown in Tab. II, we achieve the best performance for all metrics. On KITTI 2015, MonSter++ surpasses CREStereo [30] and Selective-IGEV [5] by 18.93% and 11.61% on the D1-all metric of all regions, respectively. As for KITTI 2012, we significantly outperform the existing SOTA by a large margin. Compared to the current highest-precision methods, IGEV++ [33] and NMRF-Stereo [91], we achieve a 16.26% improvement and a 20.74% improvement respectively in the Out-2 metric and the Out-3 metric across all regions.
IV-C Multi-view Stereo Benchmark Performance
To demonstrate the outstanding performance of our method in the multi-view stereo task, we evaluate MonSter++ on the two popular multi-view stereo depth estimation benchmarks: DDAD [24] and KITTI Eigen split [25].
DDAD [24]. We employ the same training and testing scheme to all methods, following [13]. For all methods, we train from scratch on DDAD training set. During testing, we evaluate all ring cameras to report the average performance, instead of using only the front-view camera. As shown in Tab. III, our MonSter++ achieves the state-of-the-art (SOTA) performance on DDAD. Compared with the previous best method [13], MonSter++ achieves over 14.8% improvement on AbsRel error.
KITTI [25]. KITTI Eigen split [25] is a classical benchmark for multi-view depth estimation. Thus, we compare our method with state-of-the-art methods on this dataset. As shown in Tab. III, MonSter++ outperforms all existing methods across all evaluation metrics. It is worth mentioning that MonSter++ achieves a 18.11% improvement and a 21.21% improvement respectively on SqRel error compared with the current highest-precision methods [15] and [13].
IV-D Performance in Ill-posed Regions